A 70 g bullet moving east 400m/s strikes a 1.3 kg pendulum block on the right and become embedded on it. How high will the block-bullet rise above its original position?


1
Expert's answer
2021-09-29T09:55:29-0400

Given:

"m=0.07\:\rm kg"

"u=400\:\rm m/s"

"M=1.3\:\rm kg"


The law of conservation of momentum gives

"mu=(m+M)v"

The law of conservation of energy says

"(m+M)v^2\/2=(m+M)gh"

So

"h=v^2/2g=(mu/(m+M))^2/2g\\=(0.07*400/(0.07+1.3))^2/(2*9.8)=21\:\rm m"